# On Task Vectors and Gradients

### v5

@@ 2.1 Dominance of The First Epoch
Motivated by the strong performance of one-epoch merging, we next examine how much each epoch actually contributes to the overall optimization. Fig. 3(a) plots the epoch-wise normalized gradient norm $\frac{\|\nabla_{\theta}^{(k)}L\|}{\sum_{k^{\prime}=1}^{K}\|\nabla_{\theta}^{(k^{\prime})}L\|}$.We observe that the first epoch contributes the largest share of total gradient norms across training. Even in datasets where this dominance is less pronounced, such as RESISC45 and DTD, we speculate that the initial epoch still largely determines the training direction. As depicted in Fig. 3(b), the gradients from the first 55 epochs maintain a high cosine similarity (>0.8>0.8) with the first epoch’s gradient. Thus, the effectiveness of TA arises from the alignment between fully trained and first-epoch task vectors, the latter being closer to true task-loss gradients. Fully trained vectors favor task-specific performance, while first-epoch vectors prioritize gradient approximation. Both yield similar multitask performance, but the single-epoch setting is more efficient.
(a) Normalized gradient norms after 55 epochs of finetuning.
@@ 2.2 Proof Sketch
Stage 11: Decomposing Single-Task Error
@@
Stage 22: Error Aggregation across Tasks
@@ 2.3 Bounding Second-Order Error Term $C(\{\ \theta_{\mathrm{MT}}^{(j)}\}\bigr)$
Assume the hypothesis of Proposition 1 holds and let $C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{h})$ be the error term defined in Equation 3. We assume the tasks are all classification tasks optimized with cross-entropy loss. We also add a few structural constraints on the network itself. Specifically the model is a depth-$L$ feed-forward network with weight matrices $W^{(1)},\dots,W^{(L)}$. For every layer, there exist positive constants $s_{\ell}$ such that $\lVert W^{(\ell)}\rVert_{2}\leq s_{\ell}$. The inputs are also bounded $\lVert x\rVert_{2}\leq M_{x}$. Finally, we require the activation functions to have bounded first and second derivatives, i. e., there are $\beta_{\phi},\gamma_{\phi}>0$ such that $\sup_{z}|\phi^{\prime}(z)|\leq\beta_{\phi}\quad\text{and}\quad\sup_{z}|\phi^{\prime\prime}(z)|\leq\gamma_{\phi}.$ Setting $\Pi=\prod_{\ell=1}^{L}s_{\ell}$, we obtain the following. (i) For general activations: $\bigl\|C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\|_{2}\;\leq\;T\,\binom{h+2}{2}\,|\alpha T-1|\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\phi}_{\max}}\,{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\phi}_{\max}},$ $\text{with }{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\phi}_{\max}}\;\leq\;2\,\gamma_{\phi}\,M_{x}^{2}\,\Pi^{2}\,\beta_{\phi}^{\,2L-2}\text{ and }{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\phi}_{\max}}\;\leq\;\sqrt{2}\,M_{x}\,\Pi\,\beta_{\phi}^{\,L-1}.\;$ (ii) For ReLU activations ($\gamma_{\phi}=0$, $\beta_{\phi}=1$): $\bigl\|C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\|_{2}\;\leq\;\frac{T}{2}\,\binom{h+2}{2}\,|\alpha T-1|\,H^{\text{ReLU}}_{max}G^{\text{ReLU}}_{max}$ $\text{with }{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\text{ReLU}}_{\max}}\;\leq\;\tfrac{1}{2}\sqrt{2}\,M_{x}^{3}\Pi^{3}\beta_{\phi}^{\,3L-3}\text{ and }{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\text{ReLU}}_{\max}}\;\leq\;\sqrt{2}\,M_{x}\,\Pi\,.$
@@
    $\|C\|_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,|\alpha T-1|\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H_{\max}^{\phi}}{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G_{\max}^{\phi}}\;$
@@ A.2 Proofs of Theorem 2
    $\bigl\lVert C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\rVert_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,\lvert\alpha T-1\rvert\,H_{\max}\,G_{\max},$
@@
    $\bigl\lVert C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\rVert_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,\lvert\alpha T-1\rvert\,\frac{1}{2}\sqrt{2}\,M_{x}^{3}\,\Pi^{3}\,\beta_{\phi}^{\,3L-3}.$
@@
$\leq\big(\sum_{t\in T}\sum_{\ell=0}^{h}\|\nabla^{2}\overline{L}_{t}(\theta^{(\ell)}_{\text{MT}})\|_{2}\big)\cdot\big(\sum_{m=0}^{\ell}\sum_{t^{\prime}\neq t,t^{\prime}\in T}\big\|\alpha\nabla\overline{L}_{t^{\prime}}(\theta^{(m)}_{\text{MT}})+(\alpha-1)\nabla\overline{L}_{t}(\theta^{(m)}_{\text{MT}})\big\|_{2}\big)$
@@
$\|C(\{\theta_{MT^{(j)}}\}_{j=1}^{h})\|_{2}$ $\leq\sum_{t\in T}\sum_{\ell=0}^{h}H_{max}\sum_{m=0}^{\ell}|\alpha(T+1)-1|G_{max}$ (24)
$=T\frac{(h+1)(h+2)}{2}(\alpha(T+1)-1)H_{max}G_{max}$ (25)
@@
$K$ is the number of classes and $P$ the number of parameters. Moreover we denote by $J_{\theta}f_{\theta}(x)\in\mathbb{R}^{K\times P}$ the Jacobian of the network with respect to the parameters, $\nabla^{2}_{\theta}f_{\theta}$ is a 33-dimensional tensor in $\mathbb{R}^{K\times P\times P}$. By chain rule we can obtain that:

### v6

@@ 2.1 Dominance of The First Epoch
Motivated by the strong performance of one-epoch merging, we next examine how much each epoch actually contributes to the overall optimization. Fig. 3(a) plots the epoch-wise normalized gradient norm $\frac{\|\nabla_{\theta}^{(k)}L\|}{\sum_{k^{\prime}=1}^{K}\|\nabla_{\theta}^{(k^{\prime})}L\|}$.We observe that the first epoch contributes the largest share of total gradient norms across training. Even in datasets where this dominance is less pronounced, such as RESISC45 and DTD, we speculate that the initial epoch still largely determines the training direction. As depicted in Fig. 3(b), the gradients from the first $5$ epochs maintain a high cosine similarity (>0.8>0.8) with the first epoch’s gradient. Thus, the effectiveness of TA arises from the alignment between fully trained and first-epoch task vectors, the latter being closer to true task-loss gradients. Fully trained vectors favor task-specific performance, while first-epoch vectors prioritize gradient approximation. Both yield similar multitask performance, but the single-epoch setting is more efficient.
(a) Normalized gradient norms after $5$ epochs of finetuning.
@@ 2.2 Proof Sketch
Stage $1$: Decomposing Single-Task Error
@@
Stage $2$: Error Aggregation across Tasks
@@ 2.3 Bounding Second-Order Error Term $C(\{\ \theta_{\mathrm{MT}}^{(j)}\}\bigr)$
Assume the hypothesis of Proposition 1 holds and let $C(\{\theta_{\text{MT}}^{(j)}\}_{j=1}^{h})$ be the error term defined in Equation 3. We assume the tasks are all classification tasks optimized with cross-entropy loss. We also add a few structural constraints on the network itself. Specifically the model is a depth-$L$ feed-forward network with weight matrices $W^{(1)},\dots,W^{(L)}$. For every layer, there exist positive constants $s_{\ell}$ such that $\lVert W^{(\ell)}\rVert_{2}\leq s_{\ell}$. The inputs are also bounded $\lVert x\rVert_{2}\leq M_{x}$. Finally, we require the activation functions to have bounded first and second derivatives, i. e., there are $\beta_{\phi},\gamma_{\phi}>0$ such that $\sup_{z}|\phi^{\prime}(z)|\leq\beta_{\phi}\quad\text{and}\quad\sup_{z}|\phi^{\prime\prime}(z)|\leq\gamma_{\phi}.$ Setting $\Pi=\prod_{\ell=1}^{L}s_{\ell}$, we obtain the following. (i) For general activations: $\bigl\|C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\|_{2}\;\leq\;T\,\binom{h+2}{2}\,|\alpha T+1|\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\phi}_{\max}}\,{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\phi}_{\max}},$ $\text{with }{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\phi}_{\max}}\;\leq\;2\,\gamma_{\phi}\,M_{x}^{2}\,\Pi^{2}\,\beta_{\phi}^{\,2L-2}\text{ and }{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\phi}_{\max}}\;\leq\;\sqrt{2}\,M_{x}\,\Pi\,\beta_{\phi}^{\,L-1}.\;$ (ii) For ReLU activations ($\gamma_{\phi}=0$, $\beta_{\phi}=1$): $\bigl\|C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\|_{2}\;\leq\;\frac{T}{2}\,\binom{h+2}{2}\,|\alpha T+1|\,H^{\text{ReLU}}_{max}G^{\text{ReLU}}_{max}$ $\text{with }{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H^{\text{ReLU}}_{\max}}\;\leq\;\tfrac{1}{2}\sqrt{2}\,M_{x}^{3}\Pi^{3}\beta_{\phi}^{\,3L-3}\text{ and }{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G^{\text{ReLU}}_{\max}}\;\leq\;\sqrt{2}\,M_{x}\,\Pi\,.$
@@
    $\|C\|_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,|\alpha T+1|\,{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}H_{\max}^{\phi}}{\color[rgb]{0.72,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.72,0,0}\pgfsys@color@cmyk@stroke{0}{0.89}{0.94}{0.28}\pgfsys@color@cmyk@fill{0}{0.89}{0.94}{0.28}G_{\max}^{\phi}}\;$
@@ A.2 Proofs of Theorem 2
    $\bigl\lVert C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\rVert_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,\lvert\alpha T+1\rvert\,H_{\max}\,G_{\max},$
@@
    $\bigl\lVert C(\{\theta^{(j)}_{\mathrm{MT}}\}_{j=1}^{h})\bigr\rVert_{2}\;\leq\;T\,\frac{(h+1)(h+2)}{2}\,\lvert\alpha T+1\rvert\,\frac{1}{2}\sqrt{2}\,M_{x}^{3}\,\Pi^{3}\,\beta_{\phi}^{\,3L-3}.$
@@
$\leq\big(\sum_{t\in T}\sum_{\ell=0}^{h}\|\nabla^{2}\overline{L}_{t}(\theta^{(\ell)}_{\text{MT}})\|_{2}\big)\cdot\big\|\sum_{m=0}^{\ell}\sum_{t^{\prime}\in T}\alpha\nabla\overline{L}_{t^{\prime}}(\theta^{(m)}_{\text{MT}})-1\nabla\overline{L}_{t}(\theta^{(m)}_{\text{MT}})\big\|_{2}$
$\leq\big(\sum_{t\in T}\sum_{\ell=0}^{h}\|\nabla^{2}\overline{L}_{t}(\theta^{(\ell)}_{\text{MT}})\|_{2}\big)\cdot\big\|\sum_{m=0}^{\ell}\sum_{t^{\prime}\in T}\alpha\nabla\overline{L}_{t^{\prime}}(\theta^{(m)}_{\text{MT}})-1\nabla\overline{L}_{t}(\theta^{(m)}_{\text{MT}})\big\|_{2}$
@@
$\|C(\{\theta_{MT^{(j)}}\}_{j=1}^{h})\|_{2}$ $\leq\sum_{t\in T}\sum_{\ell=0}^{h}H_{max}\sum_{m=0}^{\ell}|\alpha T+1|G_{max}$ (24)
$=T\frac{(h+1)(h+2)}{2}(\alpha T+1)H_{max}G_{max}$ (25)
@@
$K$ is the number of classes and $P$ the number of parameters. Moreover we denote by $J_{\theta}f_{\theta}(x)\in\mathbb{R}^{K\times P}$ the Jacobian of the network with respect to the parameters, $\nabla^{2}_{\theta}f_{\theta}$ is a $3$-dimensional tensor in $\mathbb{R}^{K\times P\times P}$. By chain rule we can obtain that: